Update: Our hosting provider indicated the connectivity issues were caused by a large scale denial-of-service attack targeting another host on the same data center. They have further informed us that they have taken the necessary measures to mitigate the attack and connectivity is now stable again. Update: Connectivity appears to have been restored. We will [...]
Update (11:38 UTC) - Data center 4 migration completed. Monitored downtime for the service was 5 minutes. Update (11:32 UTC) - Data center 1 migration completed. Monitored downtime for the service was 15 minutes. On Saturday 30th July 2011, at 11am UTC we will be performing maintenance work on data centers 1 and 4. The [...]
Update (11:30 UTC) - Data center 3 migration completed. Monitored downtime for the service was less than 3 minutes. Update (11:20 UTC) - Data center 5 migration completed. Monitored downtime for the service was 6 minutes. On Saturday 23rd July 2011, at 11am UTC we will be performing maintenance work on data centers 3 and [...]
UPDATE: the migration has completed successfully. The measured downtime for our trial users was 3 minutes. On Saturday 9th July 2011, at 11am UTC we will be migrating our trial users to a new data center. The trial service might be unavailable for up to 1 hour. However, we are expecting downtime to be substantially [...]
We have received a false alarm from Pingdom about all our data centers. However, our independent monitoring shows no issues in any of our data centers. We have informed pingdom and until they resolve the problem would keep an open eye on our running services using our internal monitoring tools. All data centers are up [...]
UPDATE: Connectivity appears to have been restored. We keep monitoring in case there are further issues. UPDATE: Our hosting provider reported wide scale networking affecting the whole data center, resulting from network backbone issues. We will be monitoring the situation and take action to move to a different data center if the situation continues for [...]
Our hosting provider has reported a Distributed Denial of Service (DDOS) attack on its infrastructure, affecting the service to our trial users. Data center 2 was unavailable for approximately 17 minutes. Service is now restored and back to normal.
UPDATE: Whilst not observing any connectivity issues since the last update, we were made aware of some further work by our hosting provider to replace some network equipment. This might cause further intermittent issues. We are carefully monitoring the situation and will take action if necessary. Our hosting provided reported that the problem was caused [...]
One of our hosting providers reported connectivity issues which may affect data center 3 and 5. We are awaiting further updates whilst investigating. Service is still available, however high latency and packet loss issues have been observed.
All scheduled maintenance tasks completed successfully. No errors or downtime were observed during the maintenance period.